OpenStudy (anonymous):

find the limit: lim((square root of 4x^2+5x)-2x)) as x approaches to negative infinity. Show the process. find the limit: lim((square root of 4x^2+5x)-2x)) as x approaches to negative infinity. Show the process. @Mathematics

OpenStudy (anonymous):

i think it is \[\frac{5}{4}\]

OpenStudy (anonymous):

oh scratch that. x is going to minus infinity. i changed my mind.

OpenStudy (anonymous):

if x is going to minus infinity it is not "indeterminate" form. it is just \[\infty+\infty=\infty\]

OpenStudy (anonymous):

trick question!

OpenStudy (agreene):

Actually, this isnt an indeterminate form, nor is it complex infty. You just need to do some re-arranging, it will come out to infinity.

OpenStudy (agreene):

\[\large\sqrt{\lim_{x \rightarrow -\infty} 4x^2+5x}-2\lim_{x\rightarrow -\infty}x\] Is the same thing as the original question.

OpenStudy (agreene):

So, we can see we get infty+infty = infty which is defined.

OpenStudy (anonymous):

so what is the process to solution? can you write more?

OpenStudy (agreene):

That basically is the process, Once you move the limits, it's clear you are going to have: infty - -infty infty+infty = infty.

OpenStudy (anonymous):

oh ok. thanks.